OpenStudy (anonymous):

Integral Problem:

OpenStudy (anonymous):

\[\int\limits_{}^{} 9x/\sqrt{1-81x^4}\]

OpenStudy (gorv):

okk take x^2=t

OpenStudy (gorv):

2x*dx=dt x8dx=dt/2

OpenStudy (anonymous):

It is multiple choice: A) (1-9x^2)^1/2 + C B) 1/2(1-9x^2)^3/2 + C C) -1/2arcsin(9x^2) + C D) arcsin(9x^2) + C E) 1/2arcsin(9x^2) + C

OpenStudy (anonymous):

Did you get t=x^2 because you divided the variable because that is the arcsin formula?

OpenStudy (anonymous):

How did you get t=x^2?

OpenStudy (gorv):

in ur numerator x is there so if we have to substitue a variable whose differentiation =x

OpenStudy (gorv):

in numerator x^4 is there x^4 diff=4x^3 x^3 diff=3x^2 x^2 difff= 2x so we will take x^2=t

OpenStudy (anonymous):

Okay maybe that's my problem. Why did you take the third derivative?

OpenStudy (gorv):

on differentiation 2x*dx=dt x*dx=dt which convert numerator x*dx into dt so we left with int(9/sqrt(1-81*t^2) * dt/2

OpenStudy (gorv):

i was just showing u wat we should take so that after differentiation we get x*dx

OpenStudy (gorv):

if we take x^4=t on diff 4x^3*dx=dt

OpenStudy (gorv):

if we take x^3=t on diff 3x^2*dx=dt

OpenStudy (anonymous):

Oh because you want to get the x out of the numerator. Alright I got you.

OpenStudy (gorv):

but we need x*dx so if we take x62=t on diff 2x*dx=dt

OpenStudy (anonymous):

dx = dt/2x then

OpenStudy (gorv):

no lolll x*dx=dt/2

OpenStudy (anonymous):

Because we have to keep the variables separated? I'm think i'm sort of following. What is next?

OpenStudy (gorv):

look if we eleminate x from numerator than we will have variable only in denominator

OpenStudy (anonymous):

Hey I was doing a little bit of research and maybe the problem is simpler than we thought Integral(1/(a^2 - x^2)^1/2 = arcsin(x/a) + C

OpenStudy (gorv):

\[\int\limits_{}^{} \frac{ 9*x*dx }{ \sqrt{1-81(x^2)^{2}} }\]

OpenStudy (anonymous):

Is this what you were talking about? a = 1 x = 9x^2

OpenStudy (gorv):

x^2=t 2xdx=dt xdx=dt/2

OpenStudy (gorv):

replace xdx by dt/2 and x^2 by t

OpenStudy (anonymous):

1/2 arcsin(9x^2) + C ?

OpenStudy (gorv):

\[\int\limits_{}^{}\frac{ 9*\frac{ dt }{ 2} }{ \sqrt{1-81(t)^{2}} }\]

OpenStudy (anonymous):

Oh because xdx = dt/2, that is where I get the positive 1/2 in the problem

OpenStudy (anonymous):

Okay thank you! The answer is E right?

OpenStudy (gorv):

\[\frac{ 9 }{ 2 }\int\limits_{}^{}\frac{ dt }{ \sqrt{1^{2}-(9t)^{2}}}\]

OpenStudy (gorv):

loll u calculate that ..i dont remember the exact formula

OpenStudy (anonymous):

The formula is arcsin(x/a) + C I looked it up So you're saying that it's arcsin(9x) + C right?

OpenStudy (anonymous):

I'm going to go with E, but thanks anyways

OpenStudy (anonymous):

Wait a minute!

OpenStudy (anonymous):

Hallelujah!

OpenStudy (anonymous):

It's D!

OpenStudy (anonymous):

Exactly what you said lol